# UrbanGraph: Physics-Informed Spatio-Temporal Dynamic Heterogeneous Graphs for Urban Microclimate Prediction

### v1

@@ E.2 Sensitivity and Computation Performance Evaluation
Figure A5: Computational performance analysis. (a) and (c) illustrate the negative correlation between inference time / peak GPU memory and the average number of non-building nodes per window. (b) and (d) show the positive linear correlation between computational costs and the average number of edges per window.
@@ Appendix F Generalization and Visualization
Table A12: Performance on other target variables.
Model
AT $0.5650\pm.1324$
WS $0.7500\pm.0176$
MRT $0.8378\pm.2005$
RH $0.5159\pm.2039$
PET $0.8492\pm.0517$
@@ F.1 Prediction results on different architectures
Figure A6 compares UrbanGraph with non-graph and homogeneous graph baselines, which represent fundamentally different approaches to spatial modeling.
@@
Figure A7 provides a comparison against generative graph models and temporal variants, assessing different graph learning strategies and sequence modeling components.
@@ F.2 Performance on multiple targets
To validate the robustness of our physics-informed graph representation and the UrbanGraph architecture, we trained the model using the same inputs for the remaining five target variables. The performance of the models for these five target variables is shown in Table A12. The R² scores for all models are above 0.5, with the performance on MRT and PET nearly matching the level achieved for UTCI, which demonstrates the effectiveness of our proposed method.
To provide a qualitative view of the model’s generalization capabilities, the following figures visualize the spatio-temporal prediction results for the five other target variables.

### v2

@@ E.2 Sensitivity and Computation Performance Evaluation
Figure A7: Computational performance analysis. (a) illustrate the negative correlation between inference time / peak GPU memory and the average number of non-building nodes per window. (b) show the positive linear correlation between computational costs and the average number of edges per window.
Scalability Test on City-Scale Grids. To verify the feasibility of deploying UrbanGraph on large-scale urban grids (as queried regarding 50k–100k nodes), we conducted a stress test on a large urban region. The region was partitioned into 36 spatial windows, comprising a total of approximately 90,000 nodes. The total edge reconstruction time for the entire 13-hour sequence was recorded at 39.56 seconds (using 16 CPU workers). This empirical evidence confirms that the computational cost for city-scale edge reconstruction is negligible compared to physics-based simulations, and the divide-and-conquer strategy effectively ensures scalability.
@@ E.4 Edge Weight Calculation
E.5 MULTI-TASK VS. SINGLE-TASK LEARNING
Table A13: Performance on target variables.
Target STL MTL
UTCI $\mathbf{0.8629\pm.0696}$ $0.7460\pm.0806$
AT $\mathbf{0.5650\pm.1324}$ $0.4080\pm.1382$
WS $\mathbf{0.7500\pm.0176}$ $0.4794\pm.0315$
MRT $\mathbf{0.8378\pm.2005}$ $0.7181\pm.1700$
RH $\mathbf{0.5159\pm.2039}$ $0.4105\pm.2067$
PET $\mathbf{0.8492\pm.0517}$ $0.6883\pm.0589$
To evaluate whether shared representations could enhance performance, we compared the proposed Single-Task Learning (STL) framework (where a separate model is trained for each variable) against a Multi-Task Learning (MTL) variant, which uses a shared UrbanGraph encoder followed by variable-specific heads. As shown in Table A13, STL consistently outperforms MTL, with the performance drop being particularly pronounced for variables with distinct physical mechanisms (e.g., Wind Speed vs. MRT). We attribute this to negative transfer arising from the inherent heterogeneity of the underlying physical processes: fluid dynamics (governing Wind Speed) and radiative transfer (governing MRT) require learning distinct and often conflicting spatial dependencies. Forcing a shared encoder to capture these diverse physical laws dilutes the specificity of the node embeddings, confirming that independent training ensures each model specializes in the specific physical operator relevant to its target.
@@ F.1 Prediction results on different architectures
Figure A9 compares UrbanGraph with non-graph and homogeneous graph baselines, which represent fundamentally different approaches to spatial modeling. Figure A10 provides a comparison against generative graph models and temporal variants, assessing different graph learning strategies and sequence modeling components.
@@ F.2 Performance on multiple targets
The robustness of our physics-informed representation and the UrbanGraph architecture is validated by the consistent performance on five remaining target variables (Table A13), where all $R^{2}$ scores exceed $0.5$ and MRT/PET performance nearly matches UTCI. Qualitative generalization capabilities for these five variables are visualized in Figures A8– A11.
@@
Appendix G Generalization on Vector Fields
G.1 Dataset Generation and Physical Setup
To evaluate the model’s generalization on vector fields, we utilized the UWF3D dataset. This dataset consists of high-fidelity CFD simulations generated using the open-source platform OpenFOAM, with the following specific configurations:
Physical Modeling. Unlike the thermal dataset generated by ENVI-met, UWF3D solves the steady-state Reynolds-Averaged Navier-Stokes (RANS) equations coupled with the standard $k-\epsilon$ turbulence model. This setup captures complex non-linear fluid dynamics such as flow separation, cavity circulation, and wake turbulence.
Geometric Configuration. The simulation domain represents a $540m\times 540m$ idealized urban block. Building geometries are randomly generated via parametric design (Rhino/Grasshopper) based on real-world prototypes, ensuring morphological diversity.
Boundary Conditions. The inflow boundary follows a logarithmic wind profile with a reference velocity of $10m/s$ at $10m$ height. The simulations operate at high Reynolds numbers ($Re\approx 10^{7}-10^{8}$), ensuring Reynolds independence consistent with real-world urban aerodynamics.
Mesh and Solver. The computational domain is discretized using an unstructured mesh containing approximately 2 to 4 million grid points per sample. This high spatial resolution allows for precise prediction of the 3D velocity vector field $\mathbf{v}=(u,v,w)$ at each spatial node.
Figure A12: Visual overview of the UWF3D dataset configuration and simulation samples. The figure illustrates the diversity of parametrically generated urban geometries and their corresponding high-fidelity CFD velocity fields. The simulations, solved via RANS equations with the $k-\epsilon$ turbulence model, explicitly capture complex aerodynamic phenomena such as flow separation and wake recirculation around varying building morphologies.
G.2 Model Setup and Comparative Implementation
To rigorously validate the framework’s adaptability, we established a benchmark between the proposed UrbanGraph and a Grid-GCN baseline. Given that this task involves purely fluid dynamics without thermal radiation or biological processes, we adapted the graph construction rules to focus exclusively on the Navier-Stokes mechanisms.
G.2.1 Baseline: Grid-GCN
The Grid-GCN serves as a control group representing the standard computer vision approach applied to physical fields.
Graph Construction. It constructs a homogeneous lattice graph on the $64\times 64$ sampling grid. Connectivity is defined strictly by Euclidean proximity using the Von Neumann neighborhood (connecting only to 4 immediate neighbors: up, down, left, right).
Mechanism. Crucially, this topology forces isotropic message passing, where features are aggregated uniformly from all spatial directions regardless of the local wind vector. This effectively treats fluid prediction as a generic image smoothing task, ignoring the directional nature of momentum transport.
G.2.2 UrbanGraph
In contrast, UrbanGraph utilizes its heterogeneous graph capabilities to encode specific fluid dynamic operators. We mapped the edge types from the main paper to this aerodynamic context:
Advection Edges. To simulate momentum transport, we construct directed edges searching upstream along the dominant wind vector. This explicitly models the advection process, allowing the network to transport flow features from upstream nodes to downstream ones (mimicking the semi-Lagrangian scheme).
Wake Edges. To capture non-local turbulence and pressure drops behind obstacles, we construct wake edges that connect building boundary nodes directly to their leeward wake zones. This is topologically analogous to the ”Shading” edges in the thermal task (both representing directional obstruction effects), enforcing physical consistency in cavity circulation regions.
G.2.3 Performance Analysis
Table A14: Performance comparison on velocity vector components ($R^{2}$).
Model u v w
Grid-GCN $0.7827\pm 0.0100$ $0.6885\pm 0.0166$ $0.6858\pm 0.0141$
UrbanGraph $\mathbf{0.8886\pm 0.0004}$ $\mathbf{0.8474\pm 0.0025}$ $\mathbf{0.7937\pm 0.0020}$
Quantitative results in Table A14show that UrbanGraph achieves an $R^{2}$ of 0.8886 for the $u$-component, significantly outperforming the Grid-GCN. The performance gap highlights a critical physical insight: Isotropic aggregation (Grid-GCN) inevitably over-smooths the vector field, failing to maintain sharp velocity gradients at separation points. Conversely, UrbanGraph’s anisotropic inductive bias successfully reconstructs complex flow patterns, such as flow separation and wake recirculation, by explicitly routing information along physically valid flow paths.